OpenStudy (anonymous):

Please, Please Help!!! I had 27 problems I need help with this one only. Can you show me how to do it step-by-step please An investment account earns 4% per year compounded annually. If the initial investment was $4,000.00, how much is in the account after 3 years? Round your answer to the nearest dollar

OpenStudy (loganator):

I can help

OpenStudy (anonymous):

great!

OpenStudy (loganator):

What I think the problem is here, would be written like. (fx) = 1.04^x X being the value in years

OpenStudy (loganator):

Wait sorry

OpenStudy (loganator):

It's actually, f(x) = 4000(1.04)^3 4000 being the y intercept and the rate of change being 4 percent anually

OpenStudy (anonymous):

ok

OpenStudy (loganator):

All you have to do to solve is plug in 3 for x, so it would be, f(3) = 4000(1.04)^3. Can you solve?

OpenStudy (anonymous):

ya

OpenStudy (loganator):

Alright, once you solve you'll have the answer you are looking for.

OpenStudy (loganator):

Also you want to leave f(3) just as it is, solve the right side of the equation.

OpenStudy (anonymous):

is this it? 4499.456

OpenStudy (loganator):

Yes, if it's a multiple choice question choose the answer closest to it.

OpenStudy (anonymous):

Ok thank you sooooo much!!!!!

OpenStudy (loganator):

Np! Medal pls? It would be appreciated.

OpenStudy (loganator):

Oh you already did, thanks!